OpenStudy (anonymous):

simplify 12/ square root of 2 . Simplify and leave in radical form.

OpenStudy (anonymous):

\[\frac{ 12 }{ \sqrt{2} }\]

OpenStudy (anonymous):

12/sq2= 12 * sq2/2=6 *sq2

OpenStudy (anonymous):

multiply top and bottom by the square root of two then cancel the two

OpenStudy (anonymous):

wait, what. neither of you guys make sense.

OpenStudy (anonymous):

answer is 6 times the square root of 2. as sattelite said.

OpenStudy (anonymous):

\[\frac{12}{\sqrt{2}}*\frac{\sqrt{2}}{\sqrt{2}}\]

OpenStudy (anonymous):



OpenStudy (anonymous):

that will give you \[\frac{12\sqrt{2}}{2}\]

OpenStudy (anonymous):

oooh, because the square root of 2 ...ohhh

OpenStudy (anonymous):

well yeah the square root of 4 is just 2

OpenStudy (anonymous):

ok, thanks ryan.

OpenStudy (anonymous):

yeah then 12 divided by 2 is just 6

OpenStudy (anonymous):

so you end up with \[6\sqrt{2}\]